OpenStudy (anonymous):

I solved this equation, but figured out I'm wrong, and can't see where I went wrong. I got x=0 x=7. Solving Quadratic Equations by Factoring 6x^2=7x

OpenStudy (phi):

6x^2 - 7x =0 factor out x x (6x-7)=0 now either x is 0 or 6x-7 is 0

OpenStudy (poopsiedoodle):

42^2=/=49. if x=7, it would be 42^2=49